OpenStudy (anonymous):

It took 16 men to dig one hole in 48 hours, how long would it take 3 men to dig one hole?

OpenStudy (anonymous):

16 men dig it in 48 hours, 3 men would take less?

OpenStudy (ksaimouli):

oops i forgot this is indirect

OpenStudy (ksaimouli):

16*48/3

OpenStudy (ksaimouli):

256

OpenStudy (anonymous):

so how long would 3 men take to dig half of the hole?

OpenStudy (ksaimouli):

dive by 2

OpenStudy (ksaimouli):

128

OpenStudy (ksaimouli):

@Cemantha this is indirect propostion u need to switch teh values up side down

OpenStudy (anonymous):

to dig half of an hole is impossible think about it

OpenStudy (anonymous):

it kinda doesnt matter either way yur gnna get 9 i learned that in pre-alg that either way yu get the same number.

hero (hero):

So Cemantha, It took 16 people 2 days to dig the hole. You mean to tell me if only 3 people dug the same hole, it would take only 9 hours?

OpenStudy (anonymous):

since everyone is questioning the work i did and have dne yu figure it out im deleting my answer .

OpenStudy (anonymous):

Cemantha, "yu" is "you" PLEASE!

OpenStudy (anonymous):

jame's wold, chill...iight

OpenStudy (anonymous):

dont correct me . i type the way i type i dnt need yu correcting me thank you . i know how to speak proper

OpenStudy (anonymous):

It hurts me, it physically hurts me. "properly" ffs

OpenStudy (anonymous):

Sorry, I will shut up.

OpenStudy (anonymous):

listen leave me and my grammer ALONE.!

hero (hero):

It is indirect proportion: xy = k (16)(48) = k 768 = k 3y = 768 y = 768/3 y = 256